# Tables as Images? Exploring the Strengths and Limitations of LLMs on Multimodal Representations of Tabular Data

### v3

@@ Abstract
In this paper, we investigate the effectiveness of various LLMs in interpreting tabular data through different prompting strategies and data formats. Our analysis extends across six benchmarks for table-related tasks such as question-answering and fact-checking. We introduce for the first time the assessment of LLMs’ performance on image-based table representations. Specifically, we compare five text-based and three image-based table representations, demonstrating the influence of representation and prompting on LLM performance. Our study provides insights into the effective use of LLMs on table-related tasks.
Tables as Images? Exploring the Strengths and Limitations of LLMs on Multimodal Representations of Tabular Data
Naihao Deng*, Zhenjie Sun*, Ruiqi He, Aman Sikka, Yulong Chen, Lin Ma, Yue Zhang, Rada Mihalcea University of Michigan Westlake University {dnaihao, zjsun}@umich.edu
@@ 1 Introduction
Recent years have witnessed an explosion of Large Language Models (LLMs), with impressive performances on various Natural Language Processing (NLP) tasks Brown et al. (2020); Touvron et al. (2023); Team et al. (2023). Research to date has examined the performance of LLMs for various aspects and abilities Bang et al. (2023b); Bubeck et al. (2023); Akter et al. (2023), but their effectiveness on structured data such as tables is less explored.
@@
In this paper, we systematically evaluate model performance on tabular data for both textual LLMs and multi-modal LLMs. Specifically, we investigate several research questions, including the effectiveness of image-based representation of tabular data and how different text-based or image-based prompt methods affect LLMs’ performance on table-related tasks. Our findings include:
@@ 2 Related Work
Researchers have investigated various ways to encode tabular data. Hwang et al. (2019); Liu et al. (2021) linearize the table content. Others employ model-specific techniques such as adapting the attention mechanism to better align transformer-based models with the tabular data Zhang et al. (2020); Yang et al. (2022) or designing hierarchical encoding to capture the table structure Wang et al. (2021), fine-tuning LLMs on tabular data Zha et al. (2023), etc. In contrast, our work focuses on exploring various table representations and prompts LLMs directly.
@@
Researchers have prompted LLMs to evaluate LLMs’ performance on traditional NLP tasks Bang et al. (2023a) as well as on various complex reasoning tasks Jin et al. (2022); Wu et al. (2023). On the contrary, to the best of our knowledge, few works have prompted these LLMs on tasks involving tabular data.
@@ 3.2 Prompting Strategies
On top of different methods to represent tables, we test the vanilla prompting, chain-of-thought prompting Wei et al. (2022), and expert prompting Xu et al. (2023a) by adding “let’s pretend you are an expert in reading and understanding tables” to the prompt. Appendix D provides an example for each table representation and prompting method.
@@ 3.3 Datasets
We make use of six previously introduced datasets that cover different table sources such as Wikipedia and financial reports, examine model abilities such as information extraction and arithmetic reasoning, and cover table-related tasks such as table question answering, table fact-checking, and table-to-text generation. Table 3 provides information for each dataset we use. Considering the limited access to LLMs’ APIs and the scale of the comparison, we randomly select 100 examples from the test set for each of these datasets to conduct our analysis.
@@ 4 Research Questions
We compare using the linearized table representation (Vanilla-T in text-based table representation) and the table image (Vanilla-V in image-based table representation) as the input for both GPT-4 model and Gemini. We use vanilla prompting in this comparison and plot Figure 3. We report the results for other prompting methods in Figures 9 and 10 in Section B.1.
@@
We compare the five text-based table representations introduced in Section 3.2. On top of the five representations, we also compare how vanilla, chain-of-thought, and expert prompting affect the model performance. We conduct the comparison using all six LLMs in Section 3.1 and average their accuracy scores across FinQA, LogicNLG, TabFact, and WikiTQ. Section B.2 reports LLMs’ performance on E2E and ToTTo datasets.
@@
We test the three image-based table representations in Section 3.2 together with vanilla, chain-of-thought, and expert prompting. We test the Gemini and GPT-4 model which can take images as the input. We average the accuracy scores across FinQA, LogicNLG, TabFact, and WikiTQ. Section B.3 reports LLMs’ performance on E2E and ToTTo datasets.
@@
We test all six LLMs in Section 3.1 on FinQA which provides relevant cells in the table for each instance. We refer to the relevant cells by adding “Please pay attention to the highlighted cells: (row index, column index, cell value)” in the text prompt, or mark them on the table image directly. Appendix D provides our prompt examples. We use vanilla prompting in this comparison.
@@ 5 Open Problems to Increase the Performance of LLMs on Tabular Data
Appendix E provides more examples from each dataset where LLMs can and cannot answer well.
@@ 8 Limitations
In this study, we do not intend to exhaust every possible text representation, image representation of tables, or every possible LLM. Moreover, we do not have access to the closed-source LLMs behind their API. We hope our findings and insights in this paper can inspire future research on table-related tasks.
@@ Appendix A Contributions
Naihao Deng drafted the paper. Zhenjie Sun drafted the prompting strategies and metrics, and actively got involved in discussions of result analysis. Naihao Deng and Zhenjie Sun plotted all the tables and figures. Naihao Deng, Zhenjie Sun selected examples that appeared in this paper. Ruiqi He, Aman Sikka provided suggestions on example selections. Rada Mihalcea, Yue Zhang, Lin Ma, Yulong Chen reviewed the paper and provided feedback.
Acknowledgements
We thank Inderjeet Nair for early discussion on different forms of table representations. We thank Alexander Hanbo Li, Patrick Ng for discussions of our work and pointers to related works. We thank the authors of the table-related datasets we use in our experiments for making their datasets publically available. We thank OpenAI for providing credits to access their APIs.
Appendix B Research Questions Cont’d
B.1 RQ1 Cont’d. Can we use image-based representations of tabular data?
B.2 RQ2 Cont’d. How do different text-based prompt methods affect LLMs’ performance on tabular-related tasks?
@@ Appendix B Research Questions Cont’d
B.3 RQ3 Cont’d. How do different image-based prompt methods affect LLMs’ performance on tabular-related tasks?
@@
Appendix C Comparison of LLaMa Models and GPT-4 Models
@@
Appendix D Prompt Examples
@@
Appendix E LLMs’ Generation Examples on Each Dataset

### v4

# Tables as Texts or Images: Evaluating the Table Reasoning Ability of LLMs and MLLMs

@@ Abstract
In this paper, we investigate the effectiveness of various LLMs in interpreting tabular data through different prompting strategies and data formats. Our analyses extend across six benchmarks for table-related tasks such as question-answering and fact-checking. We introduce for the first time the assessment of LLMs’ performance on image-based table representations. Specifically, we compare five text-based and three image-based table representations, demonstrating the role of representation and prompting on LLM performance. Our study provides insights into the effective use of LLMs on table-related tasks.
Tables as Texts or Images:
Evaluating the Table Reasoning Ability of LLMs and MLLMs
Naihao Deng*, Zhenjie Sun*, Ruiqi He, Aman Sikka, Yulong Chen, Lin Ma, Yue Zhang, Rada Mihalcea University of Michigan University of Cambridge Westlake University {dnaihao, zjsun}@umich.edu
@@ 1 Introduction
Recent years have witnessed an explosion of Large Language Models (LLMs), with impressive performance on various Natural Language Processing (NLP) tasks Brown et al. (2020); Touvron et al. (2023); Team et al. (2023). Research to date has examined the performance of LLMs for various aspects and abilities Bang et al. (2023b); Bubeck et al. (2023); Akter et al. (2023), but their effectiveness on structured data such as tables is less explored.
@@
In this paper, we systematically evaluate model performance on tabular data for both textual LLMs and multi-modal LLMs. Specifically, we investigate several research questions, including the effectiveness of image-based representation of tabular data and how different text-based or image-based prompt methods affect LLMs’ performance on table-related tasks. In addition, we provide analysis and hypothesis of LLMs’ behaviors. Our findings include:
@@ 2 Related Work
Researchers have investigated various ways to encode tabular data. Hwang et al. (2019); Liu et al. (2021); Cong et al. (2024) linearize the table content. Others employ model-specific techniques such as adapting the attention mechanism to better align transformer-based models with the tabular data Zhang et al. (2020); Yang et al. (2022) or designing hierarchical encoding to capture the table structure Wang et al. (2021), further tuning LLMs on tabular data Zha et al. (2023); Zhang et al. (2023), etc. In contrast, our work focuses on exploring various table representations and prompts LLMs directly.
@@
Researchers have prompted LLMs to evaluate LLMs’ performance on traditional NLP tasks Bang et al. (2023a) as well as on various complex reasoning tasks Jin et al. (2022); Wu et al. (2023); Zheng et al. (2024). On the contrary, to the best of our knowledge, few works have prompted these LLMs on tasks involving tabular data.
@@ 3.2 Prompting Strategies
On top of different methods to represent tables, we test the vanilla prompting, chain-of-thought prompting Wei et al. (2022), and expert prompting Xu et al. (2023a) by adding “let’s pretend you are an expert in reading and understanding tables” to the prompt. Appendix E provides an example for each table representation and prompting method.
@@ 3.3 Datasets
We make use of six previously introduced datasets that cover different table sources such as Wikipedia and financial reports, examine model abilities such as information extraction and arithmetic reasoning, and cover table-related tasks such as table question answering, table fact-checking, and table-to-text generation. Table 3 provides information for each dataset we use. Considering the limited access to LLMs’ APIs and the scale of the comparison, we randomly select 100 examples from the test set for each of these datasets to conduct our analysis. In total, we run the our experiments on 54,000 instances***5 text form table representations 3 prompting strategies 6 LLMs 6 datasets 100 examples / dataset + 3 image table representations 3 prompting strategies 2 MLLMs 6 datasets 100 examples = 54,000 instances. This does not include the number of additional experiments we run on other models such as Gemma, Llama-3 and GPT-4o.. In addition, we have observed significant differences between LLMs’ performance and report the results from three significance tests in Appendix B.
@@ 4 Research Questions
We compare using the linearized table representation (Vanilla-T in text-based table representation) and the table image (Vanilla-V in image-based table representation) as the input for both GPT-4 model and Gemini. We use vanilla prompting in this comparison and plot Figure 3. We report the results for other prompting methods in Figures 9 and 10 in Section C.1.
@@
We compare the five text-based table representations introduced in Section 3.2. On top of the five representations, we also compare how vanilla, chain-of-thought, and expert prompting affect the model performance. We conduct the comparison using all six LLMs in Section 3.1 and average their accuracy scores across FinQA, LogicNLG, TabFact, and WikiTQ. Section C.2 reports LLMs’ performance on E2E and ToTTo datasets.
@@
We test the three image-based table representations in Section 3.2 together with vanilla, chain-of-thought, and expert prompting. We test the Gemini and GPT-4 model which can take images as the input. We average the accuracy scores across FinQA, LogicNLG, TabFact, and WikiTQ. Section C.3 reports LLMs’ performance on E2E and ToTTo datasets.
@@
We test all six LLMs in Section 3.1 on FinQA which provides relevant cells in the table for each instance. We refer to the relevant cells by adding “Please pay attention to the highlighted cells: (row index, column index, cell value)” in the text prompt, or mark them on the table image directly. Appendix E provides our prompt examples. We use vanilla prompting in this comparison.
@@
RQ5 and RQ6
We include two additional research questions and our answer to them in Appendix C, including whether these LLMs can reconstruct tables from the image, and whether multimodal LLMs benefit from having both formats simultaneously as input.
@@ 5 Open Problems to Increase the Performance of LLMs on Tabular Data
Appendix F provides more examples from each dataset where LLMs can and cannot answer well.
@@ 8 Limitations
In this study, we do not intend to exhaust every possible text representation, image representation of tables, or every possible LLM. Moreover, we do not have access to the closed-source LLMs behind their API. Instead, we designed our experiments to try to explore the research questions we raised in this paper. We hope our findings and insights in this paper can inspire future research on table-related tasks.
Acknowledgements
We thank Inderjeet Nair for the early discussion on different forms of table representations; Alexander Hanbo Li, and Patrick Ng for discussions of our work and pointers to related work; the authors of the table-related datasets we use in our experiments for making their datasets publically available; the anonymous reviewers for their valuable suggestions. This project was partially supported by the OpenAI Researcher Access Program and by a Microsoft Foundational Model grant.
@@ Appendix A Contributions
Naihao Deng drafted the paper. Zhenjie Sun drafted the prompting strategies and metrics, and actively got involved in discussions of result analysis. Naihao Deng and Zhenjie Sun plotted all the tables and figures. Naihao Deng, Zhenjie Sun selected examples that appeared in this paper. Ruiqi He, Aman Sikka provided suggestions on example selections. Rada Mihalcea, Yue Zhang, Lin Ma, and Yulong Chen participated in discussions throughout the entire project and provided revisions and feedback on the paper.
Appendix B Statistical Significance Test Results
We have observed significant differences between LLMs’ performance. We conduct three statistical significance tests, including Fisher’s Exact test, McNemar’s Test, and proportion Z test for the model predictions.
With p < 0.05:
1.
For Fisher’s Exact test, we find a statistically significant difference in GPT-4 performance between T and V inputs for FinQA, LogicNLG, TabFact, WikiTQ (Figure 3), its performance between vanilla and expert prompting for table text input (Table 4), and its performance difference under vanilla and chain-of-thought prompting for image-based table representations (Table 5).
- 2.
3.
For the proportion Z test, we find a statistically significant difference in GPT-4 performance between T and V inputs for FinQA, WikiTQ (Figure 3).
Appendix C Research Questions Cont’d
C.1 RQ1 Cont’d. Can we use image-based representations of tabular data?
C.2 RQ2 Cont’d. How do different text-based prompt methods affect LLMs’ performance on tabular-related tasks?
@@ Appendix C Research Questions Cont’d
C.3 RQ3 Cont’d. How do different image-based prompt methods affect LLMs’ performance on tabular-related tasks?
@@
| GPT-4 | Gemini | Metric | |
| WikiTQ | 80.0 | 58.0 | Acc |
| TabFact | 64.0 | 61.3 | |
| LogicNLG | 48.0 | 33.3 | |
| FinQA | 61.0 | 45.0 | |
| ToTTo | 42.4 | 44.2 | ROUGE-L |
| E2E | 42.6 | 41.1 |
C.4 RQ5: Can These LLMs reconstruct tables from the image?
Test:
We conduct experiments to explore table reconstruction ideas using GPT4 on E2E and FinQA datasets.
TL;DR Answer:
Mostly yes.
Full Answer:
GPT4 successfully reconstructs 97/100 tables on the E2E dataset, and 72/100 tables on the FinQA dataset. We notice that for simpler tables, GPT4 can reconstruct tables almost perfectly. For instance, in E2E, the only “mistake” the model makes is to output “é” instead of “e” in “Rainbow Vegetarian Café”.
GPT4 also demonstrates a decent capability of reconstructing complicated tables. On FinQA, GPT-4 manages to reconstruct a 4-row10-column table with over 100 words (Table 16).
| Year Ended December 31 (in millions except rates) | 2018 | 2017 | 2016 |
| Less: Average CIB Markets Interest-Earning Assets (c) | 609,635 | 540,835 | 520,307 |
| Average Interest-Earning Assets Excluding CIB Markets | $1,619,553 | $1,639,757 | $1,581,297 |
| Net Interest Yield on Average Interest-Earning … | 2.50% (2.50%) | 2.36% (2.36%) | 2.25% (2.25%) |
| Net Interest Yield on Average CIB Markets … | 0.51 | 0.86 | 1.22 |
| Net Interest Yield on Average Interest-Earning … | 3.25% (3.25%) | 2.85% (2.85%) | 2.59% (2.59%) |
But GPT-4 is more prone to hallucination or messing up with the spatial relations when the table gets more complicated. For instance, for Table 17, GPT-4 reconstructs it as Table 18. As expected, GPT-4 fails to answer the corresponding question to this table when using a table image as the input. However, we notice that when using the text representation of the table, GPT-4 also fails to answer this question. This aligns with what we have reported in.
| Year Ended December 31 (in millions) | 2010 | 2009 | 2008 |
| U.S. | $16,568 | $6,263 | $-2,094 (2094) |
| Non-U.S. (a) | 8,291 | 9,804 | 4,867 |
| Income Before Income Tax Expense/(Benefit) and Extraordinary Gain | $24,859 | $16,067 | $2,773 |
| Year Ended December 31 (in millions) | U.S. | Non-U.S. (a) | 2010 | 2009 | 2008 |
| Income before income tax expense/(benefit)… | $16,568 | $6,263 | $-2,094 (2094) | $8,291 | $9,804 |
| $24,859 | $16,067 | $2,773 |
As expected, GPT-4 fails to answer the corresponding question to this table when using a table image as the input. However, we notice that when using the text representation of the table, GPT-4 also fails to answer this question. This aligns with what we have reported in Figure 3 that On FinQA, GPT-4 better leverages the image representations than the text representations in general.
C.5 RQ6: Do multimodal LLMs benefit from having both formats simultaneously as input?
Test:
We test GPT-4 with vanilla prompting for the six datasets.
TL;DR Answer:
Not generally true.
Full Answer:
As shown in Table 19, it is not generally true that multimodal LLMs benefit from having both formats simultaneously as input.
There are cases where passing both image and text representations would increase the performance (e.g. GPT-4 on WikiTQ and FinQA). In other cases, the performance is comparable to either passing tables as text or image representation or lies in between.
| Datasets | T+V | T | V | Metric |
| WikiTQ | 80.0 | 75.0 | 60.0 | Acc |
| TabFact | 64.0 | 71.6 | 60.5 | |
| LogicNLG | 48.0 | 51.5 | 54.1 | |
| FinQA | 61.0 | 43.0 | 57.0 | |
| ToTTo | 42.4 | 43.9 | 44.0 | ROUGE-L |
| E2E | 42.6 | 44.8 | 39.4 |
Appendix D Comparison of Llama Models and GPT-4 Models
@@
Appendix E Prompt Examples
@@
Appendix F LLMs’ Generation Examples on Each Dataset
@@
Appendix G Additional Experiments on GPT-4o, Llama-3 and Gemma models
Table 21 provides additional results for vanilla prompting for GPT-4o, Llama-3 and Gemma. We have observed that the GPT-4o performs similarly if we pass the table either through image or text.
| GPT-4o | Llama-3 | Gemma | Metric | ||||
| Vanilla-T | Vanilla-V | 8B | 70B | 2B | 7B | ||
| WikiTQ | 88.0 | 82.0 | 43.0 | 69.0 | 20.0 | 34.0 | Acc |
| TabFact | 69.7 | 70.4 | 48.5 | 71.9 | 30.2 | 41.1 | |
| LogicNLG | 54.2 | 54.6 | 25.8 | 31.1 | 15.1 | 20.1 | |
| FinQA | 71.0 | 68.0 | 44.0 | 53.0 | 2.0 | 6.0 | |
| ToTTo | 44.0 | 44.8 | 8.8 | 45.6 | 13.5 | 27.2 | ROUGE-L |
| E2E | 44.2 | 43.7 | 6.6 | 18.1 | 17.7 | 18.6 | |